Circular No. 6648

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 1997 J2 Michel Meunier reports his discovery of a new comet on May 7.9 UT,

only ∼ 6′ southwest of comet C/1997 J1, perhaps 0.5 mag brighter than C/1997 J1 but having a much slower motion. Additional observations, apparently an independent discovery, have also been reported by Philippe Dupouy and J.-F. Lahitte.

1997 UT α_{2000} δ_{2000} m_{1} Observer
May 7.91366 10^{h}54^{m}10.^{s}28 +73^{o}53′08′′.6 14.0 Meunier
7.92087 10 54 09.88 +73 53 07.6 13.9  _{”}
8.01956 10 54 04.71 +73 52 52.6 15.5 Dupouy
8.05955 10 54 02.69 +73 52 46.3 15.5  _{”}
8.08126 10 54 01.62 +73 52 44.1 15.5  _{”}
8.88024 10 53 23.25 +73 50 46.7 14.2  _{”}
8.95490 10 53 19.28 +73 50 35.1 14.2  _{”}

M. Meunier (Arbonne la Forêt). 0.20-m f/6 reflector + CCD. P. Dupouy and J.-F. Lahitte (Observatoire de Dax). 0.20-m reflector + CCD.

V1974 CYGNI P. Garnavich and J. Raymond, Harvard-Smithsonian Center for Astro-

physics, report: “Narrow-band H-alpha imaging of the nebula associated with V1974 Cyg = N Cyg 1992 (cf. IAUC 6441) shows that the region of emission is apparently expanding away from the nova. CCD images obtained with the F. L. Whipple Observatory’s 1.2-m reflector on 1996 Oct. 15 were compared with data taken by P. Challis with the same equipment on 1997 May 5. The difference between the frames shows that the brightest emission region, ∼ 2′ southwest of the nova, has increased in extent by 10′′ to the southwest. A long, faint emission arc extending south of the nova around to the east with a radius of 6′–7′ is visible at both epochs. The arc has now increased in distance from the nova by ∼ 30′′ to the southeast. Another faint arc 4′ to the west has brightened but has not significantly changed position. The shift of 30′′ over a 200-day interval is consistent with the speed of light at distance of 1 kpc, suggesting that the changes are not due to physical motion of the gas. Spectra of the nebula southwest of the nova were obtained on 1996 Oct. 14 with the Multiple-Mirror Telescope and show only narrow Balmer emission. Such a spectrum is inconsistent with either reflection or ionization as the source of the emission. Instead, Lyman photons from the nova may be reprocessed by the ambient hydrogen in the interstellar gas.”

1997 May 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Brian G. Marsden